end te o Io the 19 S children d 23f Thst the rish of eongit 0 4 ppoiste) What wowld she prote be for thi HP 36,HP 21. (4 pointe) What would be the standard esror for thie problenn ) 0.011122 b) 0.0133 e) 0.7911 d) 10- 22. (4 points) What is resulting test statistics? a) 0.89 b) 0.623 c) 0.712 d) 0.99 23. (4 points) What is the resulting p-value? a) 0.1867 b) 0.224 c) 0.712 d) 0.1494
Answer #1

Solution20:

given to test for increase in abnormatlities or not

H0:p=0.05

Ha:p>0.05

MARK OPTION A

Solution21:

standard error=sqrt(p*(-p)/n

=(0.05*(1-0.05)/384

=0.011122

MARK OPTION A

Solution22:

z=p^-p/sqrt(p*(1-p)/n

p^=23/384=0.05989583

z=0.05989583-0.05/sqrt(0.05*(1-0.05)/384)

z=0.89

MARK OPTION A

0.89

Solution23:

p=1-NORM.S.DIST(0.89;1)

p=0.1867

ANSWER:

0.1867

MARK OPTION A: